# Best Foot Forward: Robust Foot Reconstruction in-the-wild

### v1

@@ 6.1 Experimental Setup
To evaluate our foot completion module, we fit three established foot models to incomplete and completed scans. The first, a PCA-based method [1], uses functional maps [19, 17] for vertex correspondences and fits a PCA model to mesh displacement vectors. The second leverages the SUPR foot model [18], while the third, FIND [3], offers a large latent space for shape and pose control, with added transformation parameters for better alignment. For all methods, we optimize shape, pose, and transformation parameters via gradient descent, minimizing Chamfer distance with the Adam optimizer [11]. Final accuracy is assessed using Chamfer and Hausdorff distances.
We evaluate our method in an end-to-end reconstruction setting using unposed video images, benchmarking against two widely used pipelines: (1) COLMAP [23], which reconstructs 3D geometry via SfM and MVS, and (2) Gaussian Opacity Fields [27], a state-of-the-art differentiable rendering method that optimizes 3D Gaussians from image observations to generate a mesh. Using 30 consumer-captured videos with varying conditions, three expert clinicians assessed randomized renders from each baseline and our method. They rated reconstructions on a 5-point scale for (1) anatomical accuracy, (2) completeness, and (3) realism.
6.2 Experimental Results
Point Completion: Table 1 demonstrates that our point cloud-based method consistently outperforms template-based approaches on incomplete point clouds, yielding lower Chamfer and Hausdorff distances. Furthermore, meshes reconstructed through our pipeline, when meshed using SPSR, achieve the lowest Chamfer distances among all meshed, further validating our design choice. Qualitative results in Figure 3 further illustrate these improvements.
@@ 6.2 Experimental Results
Table 1: Here we present the average chamfer distance (CD) and Haussdorf distance (HD). The quoted plus/minus range refers to 1 standard deviation over the test dataset.
@@
View-Completeness: We evaluate our shape completion module under varying levels of partialness using a simulated scanning setup. By limiting the maximum angle between the camera-to-foot vector, we control foot coverage—smaller angles mean more missing data. As shown in Fig. 4, increasing this angle lowers the Chamfer Distance, improving reconstruction accuracy. Notably, the error plateaus around 90, a feasible range for a person performing self-scanning, indicating that this range balances practicality and accuracy, further validating our design choice for a robust completion module.
End-to-End Pipeline: We present our results in Fig. 5 (a)–(c). Our method consistently outperforms COLMAP and GOF across all metrics. COLMAP shows the worst anatomical accuracy with major deviations, while GOF is inconsistent. Our method achieves the highest, most consistent ratings in fidelity, completeness, and surface quality. These results confirm its suitability for clinical applications like foot orthotic design and precision insole manufacturing.
@@ 7 Discussion
Our results demonstrate that our end-to-end pipeline significantly improves foot geometry reconstruction, achieving lower Chamfer and Hausdorff distances while maintaining consistency with input data. The foot completion module, leveraging learned priors, successfully reconstructs plausible geometries from sparse data, addressing the limitations of template-based methods, which showed constrained shape variability in our evaluations. Our approach enables robust reconstruction across diverse and incomplete inputs, as reflected in its consistently high surface completeness scores. Furthermore, by integrating completion and canonicalization, our method effectively mitigates occlusions and partial view challenges, leading to more accurate and reliable reconstructions, as evidenced by its superior anatomical fidelity and quality ratings. While we have shown our model performs well across a diverse range of test cases, it does lack explicit uncertainty quantification for extreme out-of-distribution foot geometry; we will seek to address this in future work.

### v2

@@ 6.1 Experimental Setup
Table 1: Average Chamfer Distance (CD) and Hausdorff Distance (HD) results. The top section shows results from direct fitting to partial inputs, middle is results for our generated point cloud, and bottom is results from fitting on our generated point clouds. Reported values include $\pm$1 standard deviation over the test dataset.
@@
Completion Module Evaluation: To assess the efficacy of our foot completion module, we perform a quantitative comparison against three established foot modeling techniques. Each method is tasked with fitting a model to both the original incomplete scans and our completed point clouds. The baseline methods are: (1) a PCA-based model [1] with vertex correspondences obtained using functional maps [28, 24]; (2) the parametric SUPR foot model [27]; and (3) FIND [4], which utilizes a large latent space for detailed shape and pose control. For all methods, we optimize shape, pose, and transformation parameters via gradient descent with the Adam optimizer [16], minimizing the Chamfer distance. Final reconstruction accuracy is quantified using both Chamfer and Hausdorff distances.
End-to-End Pipeline Evaluation: We evaluate the performance of our entire reconstruction pipeline using video footage captured in uncontrolled, real-world settings. Our method is benchmarked against two prominent 3D reconstruction pipelines: (1) COLMAP [32], a standard Structure-from-Motion (SfM) and Multi-View Stereo (MVS) pipeline that reconstructs dense geometry, and (2) Gaussian Opacity Fields (GOF) [38], a state-of-the-art method for adaptive surface reconstruction from images. Our dataset consists of 30 videos from smartphone videos, featuring diverse subjects, lighting, and backgrounds. For qualitative assessment, three expert clinicians in foot anatomy and orthotics reviewed randomized renderings from all three methods. They rated each reconstruction on a 5-point Likert scale based on three criteria: (1) anatomical accuracy, (2) completeness of the foot structure, and (3) suitability for orthotic design.
6.2 Experimental Results
Completion Module Performance: As detailed in Table 1, our point cloud completion method significantly outperforms the template-based fitting approaches when applied to incomplete data, achieving lower Chamfer and Hausdorff distances. Furthermore, when the outputs are meshed using Screened Poisson Surface Reconstruction (SPSR), the surfaces generated from our completed point clouds exhibit the lowest Chamfer distance, affirming the benefit of our approach for foot mesh recovery. The qualitative results in Figure 4 visually corroborate these quantitative improvements, showing more plausible and complete foot geometry.
Robustness to Partial Views: We analyzed our shape completion module’s performance under varying degrees of data incompleteness. In a simulated scanning environment, we controlled the partialness of the input scan by restricting the maximum viewing angle between the virtual camera and the foot surface. As shown in Figure 5, the reconstruction error (Chamfer Distance) systematically decreases as the viewing angle—and thus the surface coverage—increases. Notably, the error rate plateaus around 90, a viewing range that is practically achievable in self-scanning scenarios. This demonstrates that our module delivers robust and accurate completions.
@@ 6.2 Experimental Results
End-to-End Clinical Assessment: The results of our end-to-end evaluation are summarized in Figure 6 (a)–(c). The clinical assessment reveals that our method consistently outperforms both baselines across all criteria. Reconstructions from COLMAP were frequently cited for poor anatomical accuracy and major geometric deviations. While GOF performed better, its results were rated as inconsistent. In contrast, our method achieved the highest and most stable scores for anatomical fidelity, completeness, and suitability for orthotics. These findings underscore our system’s potential for reliable use in clinical applications, such as the design and manufacturing of custom foot orthotics.
@@ 7 Discussion
Our results demonstrate that our end-to-end pipeline significantly improves foot geometry reconstruction, achieving lower Chamfer and Hausdorff distances while maintaining consistency with input data. The foot completion module, leveraging learned priors, successfully reconstructs plausible geometries from sparse data, addressing the limitations of template-based methods, which showed constrained shape variability in our evaluations. Our approach enables robust reconstruction across diverse and incomplete inputs, as reflected in its consistently high surface completeness scores. Furthermore, by integrating completion and canonicalization, our method effectively mitigates occlusions and partial view challenges, leading to more accurate and reliable reconstructions, as evidenced by its superior anatomical fidelity and quality ratings. While overall performance is strong, fine structures such as toes can be more challenging to reconstruct accurately, particularly in sequences where the static scene assumption is mildly violated due to subtle motion. Additionally, the current model lacks uncertainty quantification for highly out-of-distribution geometries, which we plan to explore in future work.
@@ 8 Conclusion
Acknowledgments
This work was supported by the Engineering and Physical Sciences Research Council [EP/S023917/1].